# Semantics-Empowered Space-Air-Ground-Sea Integrated Network: New Paradigm, Frameworks, and Challenges

### v1

@@ V-B2 EYTs for remote control services
In remote control applications under the support of WNCS shown in Fig. 13, command generator (i.e., the sensor) should transmit the commands through uplink channel to remote terminal (i.e., the controller) for actuation, and the command messages may be corrupted by noise and interference of highly dynamic downlink channels in the SAGSIN. In order to resist the effect of dynamic erroneous channels and yield high ultimate effectiveness, the commands generated based on the received messages are expected to be precise at each time slot, i.e., commands should maintain the stability and convergence of the controlled process.
@@
From a transmission control perspective, [191] constructs a task-oriented multi-device communication system with multi-modal sources and transmission control. Specifically, in the considered human activity recognition task, sources including room environment and human acceleration information are firstly captured by several end devices and then transmitted over wireless channel. Then, a DL-based processor generates a command of whether to empower the monitoring cameras for video transmission based on the captured sources. By controlling the dissemination process of video sources, the task-oriented system achieves both high total transmission rate and moderate human activity recognition accuracy, which implies a positive ultimate effectiveness. For multi-device control services, the authors in [192] aim at task-oriented data compression in a multi-agent scenario, conducting a joint communication and control optimization. The optimization problem is modeled as a decentralized partially observable MDP (Dec-POMDP) to minimize the distortion caused by compression (a mean absolute error (MAE) metric which is similar to MSE) under constraint of transmission rate. A state-aggregation for information compression algorithm (SAIC) is proposed as near-optimal solution for the optimization problem, yielding low MAE and thus high ultimate effectiveness of the task-oriented communications. Moreover, [193] considers scheduling problem among multi-modal users over time-varying channel, where some of the users conduct task-oriented status updating while others conduct traditional data transmission. In order to jointly optimize the ultimate effectiveness of such multi-user system, both the significance of information (i.e., AoII) for task-oriented user data and the throughput of data-oriented user data should be considered. The joint AoII-throughput optimization problem is formulated as an MDP, and solution verifies a better trade-off between the task accuracy and throughput. As a practical example of remote control services which can be deployed in the SAGSIN, [34] study both semantic-level (meaning-level) and effectiveness-level communications in remote control over unreliable channel based on dynamic feature compression. The authors consider a classic CartPole control task actuated by two agents, where one agent serves as observer which captures and encodes the CartPole status image into symbols implicitly representing semantic features, and the other agent serves as actuator which generates a binary command of “left” and “right” based on the received semantic symbols. The problem can be modeled as a POMDP, and different levels of communications can be achieved using different reward function. The authors point our that by setting the reward function as minus MSE, the communication system is semantics-oriented; by setting the reward function as the Q-function learned by the actuator, the communication system is effectiveness-oriented, since the Q-function comprehensively reflects the cost of each decision.
@@ V-B3 EYTs for other computing-intensive services
In this part, we review some recent work on task-oriented communication systems designed for other tasks that also consume considerably large amount of computing resources. By a task-oriented design, the resource consumption from communication, computing, and task actuation can get further reduced while ensuring effectiveness of tasks, satisfying the energy-efficiency need for the SAGSIN with limited computing resources. For instance, [194] studies information sharing between mixed reality (MR) devices (e.g., wearable devices) supported by task-oriented communications. The transmitter user captures and transmits the scene information from his angle of view via semantic extractor, and the receiver user receives the information and utilize a generative network to reproduce the scene from the angle of view of receiver user. Simulation results show that the transmission rate and successful transmission probability at receiver side get enhanced by task-oriented designs. Another example is [195], in which the personal saliency of users in task-oriented communication networks is considered. The authors consider UAV-based image transmission scenarios, where UAVs conduct semantic extraction for captured images to generate a semantic triplet, and personalized user decides whether to download the whole image based on the preference score of the received semantic triplet. By formulating and solving multi-user resource allocation problem, high comprehensive utility values (i.e., the match scores between personal query and received triplets) of all the users can be achieved, hence yielding high ultimate effectiveness. Moreover, an identification-task-oriented multi-user communication system is studied in [196], where cars in the IoV extract the environment semantics and transmit to the central server, and the server conducts identification task for each car based on the integrated semantics from all the cars to ensure high identification accuracy. For tasks where asking and answering occur between transceivers, [197, 198] propose multi-user DL-enabled semantic communication systems for visual question answering (VQA) tasks, and [199] proposes a context-based semantic communication system with memory for question answering tasks, which all increase correct answering ratio of the tasks. Besides, [200] proposes a semantic-triplet-based semantic communication system to complete sentiment analysis and question answering tasks, and the proposed system achieves at least 43.5% and 52% accuracy gains respectively on the two tasks.
@@ V-C Lessons Learned from This Section
This section elaborates on how a task-oriented communication system is designed by non-symmetric transceivers. Specifically, the core of such non-symmetric transceiver design is to reduce the resource consumption of source processing, data transmission and task decision. An effectiveness-related information extractor is adopted at transmitter side to release the burden of data amount, while an intelligent actuator is utilized to substitute traditional separate reconstructor-actuator design and thus reduce computing resource consumption as compared to reconstruction-based methods. We believe that the task-oriented communication is a promising idea for semantic communications in the SAGSIN for its energy efficiency and design brevity.
@@ VI-A Precise and Unified Metrics Describing Semantics
Currently, there still lack precise and unified semantics metrics that can accurately and comprehensively measure multi-dimensional significance/meaning/effectiveness of the source data. On the one hand, the current semantic metrics, especially meaning similarity metrics, can only describe the structural or learned implicit similarity between original and received messages, which is rather inaccurate in “semantic domain” description. One of the major reasons is that semantic information theory is still in its infant age. Although there has been some state-of-the-art works[204] to refine classical semantic information theory proposed by Weaver[205], semantic communication theory is naturally abstract and vague in semantic representation, since it is based on the logical probability which cannot be explicitly calculated. Therefore, scholars are forced to seek alternative metrics that describe semantics from various perspectives, which indirectly while partly define how much semantic information is received by receiver. On the other hand, the existing semantic metrics are only suitable for a certain type/modal of data, which lack unity and generality. For instance, AoI and its variants reflect only the freshness aspect of semantics, and meaning similarity metrics such as BERT and SSIM are customized for specific modal of sources. On the other hand, considering that massiveness and multi-modality are fundamental characteristics of data in the SAGSIN scenario, the metrics that lack comprehensiveness will challenge the design and optimization on such multi-modal and multi-dimensional source data, since one-dimensional metrics are not suitable for describing the significance/meaning/effectiveness of data of other modals. Therefore, it is imperative to seek a class of precise and unified semantic metrics for semantic communication system design of multi-modal and multi-dimensional data. Our previous works [43, 206] initially make some efforts on unifying the significance metrics to propose the GoT metric as a comprehensive goal-oriented semantic metric, which describes the significance by involving the effect of all the natural cost, actuation cost, and the potential cost caused by actuation. In order to propose comprehensive meaning similarity or effectiveness metrics, some problems should be further solved. For instance, a distance-based metric in a certain domain can be considered to unify current meaning-related metrics that respectively reflect the similarity of image, text, and speech sources. Moreover, to measure the effectiveness yielded by source data (or sampled data), we may quantify the potential cost in the life cycle of a certain message, including cost of total transmission delay, cost of total energy consumption, and cost reduced by a precise actuation (or cost increased by a wrong actuation), and a weighted “unified effectiveness metric” can be finely designed.
@@ VI-B Semantic Extraction of Multi-modal Data
Semantic extraction is always the initial and essential step in DL-based semantic communications, since the performance of extractor directly affects the performance of semantics transmission and further source reconstruction. Nevertheless, current research focuses on semantic extraction for only specific modal of source data, which cannot directly extract the multi-dimensional semantics from multi-modal data widely distributed in the SAGSIN. In fact, scholars have recently discussed the semantic extraction for multi-modal image-text-fused source data in VQA task[197], while they separately extract the image semantics and text semantics by ResNet and LSTM network, respectively. This separate extractor design for different modal of sources must adopt two or more exclusive NNs, and training such semantic extractor will consume considerably large computing resources, which may not be feasible for edge devices and servers that have access to limited computing resources in the SAGSIN. Therefore, the issue of extracting semantics of massive multi-modal data with only one intelligent NN has to be addressed. For instance, videos are a typical type of multi-modal data with both image and audio sources. Instead of separating audio information from image frames and extracting the semantics respectively, we may consider image frames and audio information at the same time as semantic-related. We can cut audio signal frame by frame, and utilize state-of-the-art DL techniques such as Transformer to simultaneously extract semantic information of both audio signal and image frame in the current time slot. Because of semantic-related audio and image sources, the parameters of NNs will get reduced as compared to separate design, for semantic representation includes the relation of image and audio of current frame, and thus representation length for current frame is shorter. By integrated extractor design for multi-modal data, a lite semantic coder can be realized and may be directly trained and deployed on small IoT devices in the SAGSIN.
VI-C Heterogeneous Task-oriented Communication Networks Based on Edge Intelligence
In the SAGSIN, massive and heterogeneous tasks are actuated in a distributed manner by edge devices. Especially, even for the same goal, several cooperative devices may execute different specific intelligent tasks. In the current literature, task-oriented communication network design usually consider homogeneous task executed cooperatively by edge devices[207, 208], which cannot be directly applied for heterogeneous tasks scenario, since the former design consider each device as achieving same goal. The only exceptions are [162, 209], where both semantic communication users and technical communication users work to achieve heterogeneous goals (i.e., both semantic and bit-level reconstruction). However, in the practical applications, tasks are more complex and different users may require different level of quality of service (QoS) or quality of experience (QoE) in task actuation, which calls for more intelligent design of heterogeneous task-oriented communication networks based on edge intelligence. We may consider a joint detection-classification heterogeneous task which is executed by multiple edge devices distributed in the SAGSIN. Specifically, some of devices detect the objects, while the others classify them based on both their own perception images and detection results by detecting devices. Since the classifier input is multi-modal, i.e., consisting of both detection results and perception images, we have two methods for classifier implementation. Firstly, we can separate the classifier into perception-only and detection-based-only devices, respectively processing the perception images and detection results. Another way is that we assume that all the classifier devices receive both inputs to execute classification tasks. In order to jointly train these detection and classification devices, federated edge learning can be utilized as a solution. However, considering that the goals of devices are different, we can firstly conduct clustered training for devices with the same task or goal in the first training phase, and secondly fine-tune the parameter among devices with different goals in the second training phase. By such two-phase training, heterogeneous tasks can be actuated cooperatively to yield high ultimate effectiveness.
VI-D Joint Decision-Execution Design for the Actuator
@@
In the SAGSIN, task execution at the terminals will directly affect the environment statuses, which is the direct effectiveness yielded by the decision commands generated from actuator. However, the current execution modules (or techniques) are principally assumed as separated from the decision-maker design, which is still way from the envisioned PCCAIP in Fig. 3 of Section II. For instance, the current task-oriented image classification decision-maker only gives the category results on a certain image without considering how the classification results facilitate task execution to affect the external environment. The only exception is [34], where the intelligent actuator both gives decision results (i.e., left or right) and conducts task execution to complete CartPole game. The real tasks facilitated by the SAGSIN is far more complex and heterogeneous than simple CartPole control, which necessitates more intelligent and comprehensive techniques of joint decision-execution realization for the actuator in order to directly yield ultimate effectiveness. Here we study an abnormal status observation and rescue task to demonstrate the advantage of such joint decision-actuation design. The observers distributed in the network monitor the environment, and when detecting abnormal statuses, the observer will transmit them using SPTs, METs, or EYTs. The terminals receive the statuses using a joint reconstruction & task-decision & task-execution technique to directly conduct rescue task, e.g., sending rescuers to dangerous areas for sinking ships. The state-of-the-art works reviewed in Section V mainly focus on the monitoring steps of this task in order to enhance accuracy of the observation. However, in joint decision-execution design, we are not only interested in the accuracy of observation but also the affect of the decision (which is based on observation results) caused by task execution, e.g., recovery of abnormal statuses due to timely rescue. When several observed objects are detected as abnormal at the same time but the terminal can only support one of the object to be rescued, the actuator should be finely designed to find out the decision yielding the most ultimate effectiveness via considering the different level of significance of all the objects. In such cases, we must resort to joint design instead of only enhancing the accuracy of observation.

### v3

@@ V-B2 EYTs for remote control services
In remote control applications under the support of WNCS shown in Fig. 12, the sensor should transmit the real-time status messages through uplink channel to the controller for command generation, and the commands will be transmitted back to the operator through downlink channel. The controlling process may be corrupted by noise and interference of highly dynamic channels in the SAGSIN. In order to resist the effect of dynamic erroneous channels and yield high ultimate effectiveness, the commands generated based on the received messages are expected to be precise at each time slot, i.e., commands should maintain the stability and convergence of the controlled process.
From a transmission control perspective, [204] constructs a task-oriented multi-device communication system with multi-modal sources and transmission control. Specifically, in the considered human activity recognition task, sources including room environment and human acceleration information are firstly captured by several end devices and then transmitted over wireless channel. Then, a DL-based processor generates commands of whether to empower the monitoring cameras for video transmission based on the captured sources. By controlling the dissemination process of video sources, the task-oriented system achieves both high total transmission rate and moderate human activity recognition accuracy, which implies a positive ultimate effectiveness.
@@
For multi-device control services, the authors in [205] aim at task-oriented data compression in a multi-agent scenario, conducting a joint communication and control optimization. The optimization problem is modeled as a decentralized partially observable MDP (Dec-POMDP) to minimize the distortion caused by compression (a mean absolute error (MAE) metric which is similar to MSE) under constraint of transmission rate. A state-aggregation for information compression algorithm is proposed as near-optimal solution for the optimization problem, yielding low MAE and thus high ultimate effectiveness of the task-oriented communications. Moreover, [206] considers scheduling problem among multi-modal users over time-varying channel, where some of the users conduct task-oriented status updating while others conduct traditional data transmission. In order to jointly optimize the ultimate effectiveness of such multi-user system with multi-modal tasks (i.e., both data reconstruction and status updating tasks), both the significance of information (i.e., AoII) for task-oriented user data and the throughput of data-oriented user data should be considered. The joint AoII-throughput optimization problem is formulated as an MDP, and solution verifies a better trade-off between the status prediction accuracy of status updating tasks and the throughput of data recovery tasks.
As a practical example of remote control services which can be deployed in the SAGSIN, [44] studies both semantic-level (meaning-level) and effectiveness-level communications in remote control over unreliable channel based on dynamic feature compression. The authors consider a classic CartPole control task actuated by two agents, where one agent serves as observer which captures and encodes the CartPole status image into symbols implicitly representing semantic features, and the other agent serves as actuator which generates a binary command of “left” and “right” based on the received semantic symbols. The problem can be modeled as a POMDP, and different levels of communications can be achieved using different reward functions. The authors point out that by setting the reward function as minus MSE, the communication system is semantics-oriented; by setting the reward function as the Q-function learned by the actuator, the communication system is effectiveness-oriented, since the Q-function comprehensively reflects the cost of each decision.
@@ V-B3 EYTs for other computing-intensive services
In this part, we review some recent work on task-oriented communication systems designed for other tasks that also consume considerably large amount of computing resources. By a task-oriented design, the resource consumption from communication, computing, and task actuation can get further reduced while ensuring effectiveness of tasks, satisfying the energy-efficiency need for the SAGSIN with limited computing resources.
For instance, [207] studies information sharing between mixed reality (MR) devices (e.g., wearable devices) supported by task-oriented communications. The transmitter user captures and transmits the scene information from his angle of view via semantic extractor, and the receiver user receives the information and utilizes a generative network to reproduce the scene from the angle of view of receiver user. Simulation results show that the transmission rate and successful transmission probability at the receiver side get enhanced by task-oriented designs. Another example is [208], in which the personal saliency of users in task-oriented communication networks is considered. UAVs conduct semantic extraction for captured images to generate a semantic triplet, and personalized users decide whether to download the whole image based on the preference score of the received semantic triplet. By formulating and solving multi-user resource allocation problem, high comprehensive utility values (i.e., the match scores between personal query and received triplets) of all the users can be achieved, hence yielding high ultimate effectiveness. Moreover, an identification-task-oriented multi-user communication system is studied in [209]. Cars in the IoV extract the environment semantics and transmit to the central server, and the server conducts identification task for each car based on the integrated semantics from all the cars to ensure high identification accuracy. There are also some efforts on task-oriented communications for question asking and answering tasks. [210, 211] propose multi-user DL-enabled semantic communication systems for visual question answering (VQA) tasks, both of which increase correct answering ratio of the tasks. [212] also proposes a context-based semantic communication system with memory to enhance the correction ratio of question answering tasks. Besides, [213] proposes a semantic-triplet-based semantic communication system to complete sentiment analysis and question answering tasks, and the proposed system achieves at least 43.5% and 52% accuracy gains respectively on the two tasks.
@@ V-C Lessons Learned from This Section
This section elaborates on how a task-oriented communication system is designed by non-symmetric transceivers. Specifically, the core of such non-symmetric transceiver design is to reduce the resource consumption of source processing, data transmission and task decision. An effectiveness-related information extractor is adopted at transmitter side to release the burden of data amount, while an intelligent actuator is utilized to substitute traditional separate reconstructor-actuator design and thus reduce computing resource consumption as compared to reconstruction-based methods. We believe that the task-oriented communication is a promising idea for semantic communications in the SAGSIN for its energy efficiency and brief design.
@@ VI-A Precise and Unified Metrics Describing Semantics
Currently, there still lack precise and unified semantic metrics that can accurately and comprehensively measure multi-dimensional significance/meaning/effectiveness of the source data. On the one hand, the current semantic metrics, especially meaning similarity metrics, can only describe the structural or learned implicit similarity between original and received messages, which is rather inaccurate in “semantic domain” description. One of the major reasons is that semantic information theory is still in its infant age. Although there have been some state-of-the-art works[217] to refine classical semantic information theory[218], semantic communication theory is naturally abstract and vague in semantic representation, since it is based on the logical probability which cannot be explicitly calculated. Therefore, scholars are forced to seek alternative metrics that describe semantics from various perspectives, which indirectly while partly define how much semantic information is received by the receiver.
On the other hand, the existing semantic metrics are only suitable for a certain type/modal of data, which lack unity and generality. For instance, AoI and nonlinear AoI reflect only the freshness aspect of semantics, and meaning similarity metrics such as BERT and SSIM are customized for specific modal of sources. Also, it is an indisputable fact that massiveness and multi-modality are fundamental characteristics of data in SAGSIN scenario. The existing one-dimensional metrics are not sufficient for describing the significance/meaning/effectiveness of multi-modal and multi-dimensional source data. Therefore, it is imperative to seek a class of precise and unified semantic metrics for semantic communication system design of multi-modal and multi-dimensional data. Our previous works [55, 219] initially make some efforts on unifying the significance metrics to propose the GoT metric as a comprehensive goal-oriented semantic metric, which describes the significance by involving the effect of all the natural cost, actuation cost, and the potential cost caused by actuation. In order to propose comprehensive meaning similarity or effectiveness metrics, some problems should be further solved. For instance, a distance-based metric in a certain domain can be considered to unify current meaning-related metrics that respectively reflect the similarity of image, text, and speech sources. Moreover, to measure the effectiveness yielded by source data, we may quantify the potential cost in the life cycle of a certain message, including cost of total transmission delay, cost of total energy consumption, and cost reduced by a precise actuation (or cost increased by a wrong actuation), and a weighted “unified effectiveness metric” can be finely designed.
@@ VI-B Semantic Extraction of Multi-modal Data
Semantic extraction is always the initial and essential step in DL-based semantic communications, since the performance of extractor directly affects the performance of semantic transmission and further source reconstruction. Nevertheless, current research focuses on semantic extraction for only one specific modal of source data, which cannot directly extract the multi-dimensional semantics from multi-modal data widely distributed in the SAGSIN. In fact, scholars have recently discussed the semantic extraction for multi-modal image-text-fused source data in VQA task[210], while they separately extract the image semantics and text semantics by ResNet and LSTM network, respectively. This separate extractor design for different modals of sources must adopt two or more exclusive NNs. Training such semantic extractor will consume considerably large computing resources, which may not be feasible for end devices and edge servers that have access to limited computing resources in the SAGSIN. Therefore, the issue of extracting semantics of massive multi-modal data with only one intelligent NN has to be addressed.
For instance, videos are a typical type of multi-modal data with both image and audio sources. Instead of separating audio information from image frames and extracting the semantics respectively, we may consider image frames and audio information at the same time as semantic-related. We can cut audio signal frame by frame, and utilize state-of-the-art DL techniques such as Transformer to simultaneously extract semantic information of both audio signal and image frame in the current time slot. Because of semantic-related audio and image sources, the parameters of NNs will get reduced as compared to separate design, for semantic representation includes the relation of image and audio of current frame, and thus representation length the for current frame is shorter. By integrated extractor design for multi-modal data, a lite semantic coder can be realized and may be directly trained and deployed on small IoT devices in the SAGSIN. This promising idea has been initially explored by [220], where a joint image-audio generative diffusion model is utilized for video reconstruction. In their work, the authors quantitatively draw the conclusion that the computational complexity of the proposed model gets significantly decreased compared to baseline models. Such GAN-based model can be potentially utilized for end-to-end semantic communication systems that have smaller parameter sizes.
VI-C Joint Decision-Execution Design for the Actuator
@@
In the SAGSIN, task execution at the terminals will directly affect the environment statuses, which is the direct effectiveness yielded by the decision commands generated from actuator. However, the current execution modules (or techniques) are principally assumed as separated from the decision-maker design, which is still way from the envisioned PCCAIP in Fig. 3 of Section II. For instance, the current task-oriented image classification decision-maker only gives the category results on a certain image without considering how the classification results facilitate task execution to affect the external environment. The only exception is [44], where the intelligent actuator both gives decision results (i.e., left or right) and conducts task execution to complete CartPole game. The real tasks facilitated by the SAGSIN are far more complex and heterogeneous than simple CartPole control, which necessitates more intelligent and comprehensive techniques of joint decision-execution realization for the actuator in order to directly yield ultimate effectiveness.
Here we study an abnormal status observation and rescue task to demonstrate the advantage of such joint decision-actuation design. The observers distributed in the network monitor the environment, and when detecting abnormal statuses, the observer will transmit them using SPTs, METs, or EYTs. The terminals receive the statuses using a joint reconstruction & task-decision & task-execution technique to directly conduct rescue task, e.g., sending rescuers to dangerous areas for sinking ships. The state-of-the-art works reviewed in Section V mainly focus on the (remote) monitoring steps of this task in order to enhance the accuracy of observation. However, in joint decision-execution design, we are not only interested in the accuracy of observation but also the effect of the decision (which is based on observation results) caused by task execution, e.g., recovery of abnormal statuses due to timely rescue. When several observed objects are detected as abnormal at the same time but the terminal can only support one of the object to be rescued, the actuator should be finely designed to find out the decision yielding the most ultimate effectiveness via considering the different levels of significance of all the objects. In such cases, we must resort to joint design instead of only enhancing the accuracy of observation.
VI-D Heterogeneous Task-oriented Communication Networks Based on Edge Intelligence
In SAGSIN, massive tasks are actuated by heterogeneous nodes in a distributed manner by edge devices. Given the diverse nature of tasks, the relations among nodes may have two principal manners. For nodes which aim at the same task, they should cooperate with each other for the identical task. Conversely, for nodes which target different tasks, they may compete with each other to preempt for limited resources. Therefore, in the task-oriented communication network, it is of vital importance to study intelligent multi-dimensional resource scheduling policies among heterogeneous nodes based on edge intelligence. In the current literature, task-oriented communication network design usually consider homogeneous task executed cooperatively by edge devices[221, 222], which cannot be directly applied for heterogeneous tasks scenario, since the former design considers each device as achieving the same goal. The only exceptions are [175, 223], where both semantic communication users and technical communication users work to achieve heterogeneous goals (i.e., both semantic and bit-level reconstruction). However, in the practical applications, tasks are more complex and different users may have distinct perception qualities and computing resources, which calls for more intelligent and comprehensive design for each users based on edge intelligence.
For nodes which cooperate with each other for the same task, we may consider a multi-UAV tracking scenario where several heterogeneous UAVs track one moving object. Because of limited device capabilities of each UAV, the computing resources and perceived information are insufficient for each UAV to independently maintain the tracking task. As a result, they should resort to remote edge/cloud servers that can aggregate the perceived information from all the nodes to generate global status information (e.g. the absolute position and speed of the object). Next, the remote servers can compute moderate actuation commands (e.g. the speed and direction) for each UAV. Also, each UAV can compute their own actuation command based on local perception of the object. In order to better track the object while avoiding collision between UAVs, command merging algorithm can be utilized at each UAV to integrate the commands from both itself and the remote server. In this case, both centralized and decentralized policies are adopted, which rely highly on the proposed PCCAIP in this survey.
However, for nodes which compete with each other due to the pursuit of distinct tasks, it is worth noting that the shared communication/actuation resources must be moderately allocated to ensure high effectiveness for each task. Since the nodes serve different tasks, the feasibility of implementing centralized policies becomes physically constrained. In this case, a decentralized resource scheduling policy based on game theory can be adopted, such that the nodes can reach equilibrium in utilizing shared resources. We may consider NOMA-based multi-UAV tracking scenario where each UAV tracks its own object and communicate with a remote server using the same spectrum. UAVs must adjust their own transmitting power to mitigate interference among themselves, thereby facilitating the reliable transmission of status information to the remote server. This in turn enables the remote server to make more effective task decision for each UAV. Here, decentralized gaming among UAVs also relies highly on the proposed PCCAIP in the manuscript because of limited device capabilities in SAGSIN.